OpenStudy (anonymous):

What is the value of the expression? 4^3 • 4^−5

OpenStudy (sleepyhead314):

for something like this... x^a * x^b = x^(a+b) so 4^3 • 4^−5 becomes 4^(3+-5) = 4^(3-5) = ?

OpenStudy (anonymous):

thank you